CD was away for a conference and the weekend, but trained Michael G on collector maintenance prior to leaving. MG and MHalbur took care of Fog collections during this period, as Dave’s knee is just beginning to heal.
8/14 – no fog, collector did not open. 8/15 sample present in morning (sample was cloudy) and post clean blank was taken by MG. The hose was cleaned so should not be a contamination issue with future samples, but could be an issue in the 8/14 Fog sample. 8/16 a fog sample was retrieved by MG at 6am followed by a post clean blank the next day (8/17 at 9:00 in the morning).
On 8/18 CD received the 8/18 fog sample from MG at 9:00 am, and a post clean blank was taken in the afternoon.
**It came to our attention today that our previous samples may be slightly low on HCl additions due to an underestimation of sample volumes. Previous samples were estimated to be only 80% of there actual volume due to the use of a marked 200mL bottle when determining approx volume in 250mL sample bottles.
CD cleaned the collector on 8/12, at which time 2 strands were lost on one of the collector stand frames. The line broke, then broke again during repair, causing a small loss of two wraps around the frame. At 4pm on on 8/12 the RH was 66 and Temp 22 which was consistent with the met station.
On 8/13 there was a full sample at 9:00 am and the collector was cleaned and set up for CD’s absence in the afternoon. At 4:45 pm RH was 57 and Temp 25 which was consistent with the met station. I noticed that the hose was still a little grimy after vigorous cleaning.
Conditions were foggy upon CD’s arrival this morning, and the collector was cleaned in attempt to get a morning sample. No luck, but the RH was 95% and Temp 15.5 at 9:30 am which was consistent with the Met station.
There was a big Arctiid moth (aka tiger moth) snuzzled up on the collectors stands prior to cleaning. See pictures below! It most likely got in there while the collector opened, but failed to condense enough fog for a sample on the night prior to 8/8.
Dry, low humidity. No news except that the fog project was featured on a local NPR radio special.
A very small ~10mL sample was retrieved on 8/7 and the collector had been cleaned on 8/6. The air was clear at 9:20 in the morning on 8/7, despite heavy smoke the prior afternoon and presumably during part of the night (possible contaminant?). Time did not allow for a full collector clean on 8/7 (particularly due to low sample volume and questionable forecast). A blank was collected on 8/7 after a ‘quick-clean’ was done by only rinsing the strands in place with RO water prior. No full wash completed.
The next morning, on 8/08 the sample was too small to keep.
Torregrosa’s passive collector ( fog harp) had no fog condensation from 8/5-8/8 despite small amounts of drizzle (0.1″) during this period.
Collector cleaned and set up, and post clean blank collected today by CD. Having difficulty cleaning grime from sample tube without proper brush. The air was very smokey all day and may effect sample chemistry. At 9:30 am the RH was 47 and temp 31 which are both consistent with the met station.
Weekend and busy period! CD unable to attend to collector and there was potentially some missed fog during this time. Dave Andersen was/is out of commission due to a knee injury.
Waiting for fog after 7/31 cleaning. Collector did was not open on 8/1 am due to low overnight RH on 7/31-8/1.
CD retrieved a full fog sample at 2:00pm on 8/2 (after the morning volunteer workday). Humid conditions persisted through the morning and the sample was cool upon retrieval in afternoon. The collector was last cleaned on 7/31. CD was unable to obtain a post-clean blank for this sample until 8/6, but collector was closed and shut down between sample and blank collection period.
CD retrieved a small sample on 7/31at 9:30 am, followed by a blank after cleaning at 11:00 am. The collector had been cleaned prior on 7-30. At 5pm on 7/30 the RH was 34% and Temp 34. The both were consistent with the adjacent Met station.
The collector was shut down from 7/23-7/30. Light fog may have been missed on 7/28 and 7/29.
Pepperwood is still experiencing high humidity levels with out the presence of Fog, causing the collector to open much earlier than necessary and suck in living and non living contaminants.